Consider a 0.375 M aqueous solution of sodium cyanide (NaCN).

A) Write the reaction of the cyanide anion (CN-) with water.

B) Calculate [OH-], pOH, pH. Also, calculate percent ionization. (Obeys 5% rule). (Ka of HCN = 4.9 x 10^-10)
Answer #1

0) Reaction of CN- with Ho, CN- +40 2 HIN109) + OH lag) 6) Given, [No (N] = 0.375 M. NOIN (ag) Not + (N- 10.375 00 C -0.375m

pon = -log [or] pOH = -log (2.76X103) IPOH = 2.56 pH = 14-2156 = 11.44 I PH = 11.44 c) % ionisation = [OH I xloo [EN] 2.76770